OpenStudy (anonymous):

a company offers a starting yearly salary $30,000 with raises of $3,500 per year. Find total salary over a ten year period

OpenStudy (anonymous):

would you times 3500*9 then add that to 30000

OpenStudy (anonymous):

If the starting salary is 30,000 and it's increasing $3500 per year at ten years it will equal 65,000. 3500 * 10 years + 30,000

OpenStudy (anonymous):

it would be 3500 *9 because she does not get it the first year right

OpenStudy (anonymous):

Yes, my bad....I gave a 3500 bonus year..d'oh!

OpenStudy (anonymous):

I want that lol

OpenStudy (anonymous):

So, 3500*9+30,000

OpenStudy (anonymous):

so that means that it would be 61,500?

OpenStudy (anonymous):

mathstudent55 I am down to my last three question to have my aa yeah

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

mathstudent you still here where did you go

OpenStudy (mathstudent55):

The sum of the the terms in an arithmetic series up to term n is the average of the first and nth term multiplied by n. The first term is 30,000, the second therm is 33,500, the third term is 37,000. The nth term is an = a1 + (n -1)d, where a1 = first term, n is number of terms, and d is the difference between terms. Here you have a10 = 30,000 + (10 -1)3500 = 61,500 The sum of terms 1 thorugh 10, or 30,000 + 33,500 + ... + 61500 = 10(30,000 + 61,500)/2 = 457,500

OpenStudy (mathstudent55):

$61,500 is the salary in the 10th year. The question is how much was the total salary over the ten years. For that you need to add the salary of all years.

OpenStudy (anonymous):

I knew that and that is where I got it wrong last time dangit. oh well I guess one wrong is not going to hurt me but crapparoo

OpenStudy (anonymous):

That's where it's at....

OpenStudy (anonymous):

I'm talking about mathstudent's answer....

OpenStudy (anonymous):

can you help me with the last question I just posted

OpenStudy (mathstudent55):

i'll look for it